### v1

# RAGGED: Towards Informed Design of Retrieval Augmented Generation Systems

Abstract
Retrieval-augmented generation (RAG) greatly benefits language models (LMs) by providing additional context for tasks such as document-based question answering (DBQA). Despite its potential, the power of RAG is highly dependent on its configuration, raising the question: What is the optimal RAG configuration? To answer this, we introduce the RAGGED framework to analyze and optimize RAG systems. On a set of representative DBQA tasks, we study two classic sparse and dense retrievers, and four top-performing LMs in encoder-decoder and decoder-only architectures. Through RAGGED, we uncover that different models suit substantially varied RAG setups. While encoder-decoder models monotonically improve with more documents, we find decoder-only models can only effectively use documents, despite often having a longer context window. RAGGED offers further insights into LMs’ context utilization habits, where we find that encoder-decoder models rely more on contexts and are thus more sensitive to retrieval quality, while decoder-only models tend to rely on knowledge memorized during training. 111Code/data for the RAGGED framework are available at https://github.com/neulab/ragged
[…]
RAGGED: Towards Informed Design of
Retrieval Augmented Generation Systems
Jennifer Hsia††thanks: Equal contribution. Afreen Shaikh Zhiruo Wang Graham Neubig Carnegie Mellon University {jhsia2,afreens,zhiruow,gneubig}@cs.cmu.edu
1 Introduction
Retrieval-augmented generation (RAG) (Chen et al., 2017a; Lewis et al., 2020) is a technique widely applied to enhance the performance of top-performing LMs on knowledge-intensive generation tasks like document-based question answering (Karpukhin et al., 2020). Given a question, the technique includes using a retriever model to obtain multiple relevant passages (i.e. paragraphs) across potentially different documents, then inputting these passages to a reader model as additional contexts for generating an answer.
However, using RAG effectively is not straightforward, and existing literature provides mixed, even contradictory, suggestions for configuring RAG. While different LMs have different context length limits (Chung et al., 2022; Tay et al., 2023; Touvron et al., 2023b; Liu et al., 2023), these limits alone do not simply determine the optimal number of passages to provide in context. Early works suggest that providing more retrieved passages results in strictly better outputs (Izacard and Grave, 2021). In contrast, other works find that providing a selected set of passages (Asai et al., 2022), sentences (Wang et al., 2023), or tokens (Berchansky et al., 2023) outperforms providing the full set, presumably due to increased context relevance and quality. In addition to context quality and quantity, how robust a reader model is to noisy context also determines downstream performance. For example, Hoffmann et al. (2022) train models on high-quality data, while Yu et al. (2023) intentionally inject noisy content into input context during model training, and observe increased robustness of these models to low-quality contexts.
To provide more concrete suggestions of the best practices under various cases, we introduce an analysis framework, RAGGED,222For “retrieval augmented generation generalized evaluation device”. to test RAG combinations on a suite of representative document-based question answering (DBQA) tasks, including open-domain datasets like Natural Questions (Kwiatkowski et al., 2019) and HotpotQA (Yang et al., 2018), which respectively focus on single-hop and multi-hop questions, as well as BioASQ, which targets the specialized, biomedical domain. To ensure a comprehensive evaluation, we incorporate both classic sparse and dense retrievers, BM25 (Robertson et al., 2009) and ColBERT (Khattab and Zaharia, 2020), and four top-performing reader models in encoder-decoder and decoder-only architectures, from Flan (Chung et al., 2022; Tay et al., 2023) and LLaMa (Touvron et al., 2023b) families, respectively.
We begin by exploring ”How many contexts can readers benefit from?” (§5). Our analysis identifies the optimal context quantity for individual LMs. We find that encoder-decoder models can effectively utilize up to 30 passages within their 2-token limit, whereas decoder-only models’ performance declines beyond 5 passages, despite having twice the size of the context limit (4).
Given this intriguing difference, we investigate the models’ context utilization behaviors (§6) and ask ”How reliant are models on provided contexts?”. We find that decoder-only models, which memorize more during training, exhibit comparatively less reliance on additional, test-time contexts. In contrast, encoder-decoder models, which memorize less during training, are more faithful to the provided contexts. This suggests that providing passages for context-reliant encoder-decoder models is beneficial, whereas it is less so for memory-reliant decoder-only models.
Given that some models are reliant on context, we also examine the importance of context quality by asking “How does the retriever quality affect readers’ contextualization behavior?” (§7) Our analysis considers two aspects: a retriever’s ability to identify high-quality passages and a reader’s response to varying passage quality. While dense, neural retrievers perform better on open-domain questions, sparse, lexical retrievers readily achieve comparable accuracy on special domains, with much less computation. Neural retrievers’ advantage readily benefits encoder-decoder models, especially for single-hop questions. However, the benefits are much less pronounced for decoder-only models and multi-hop questions.
In summary, we demonstrate how RAGGED enables us to derive actionable insights about the conditions under which state-of-the-art RAG components combine to excel. We introduce a reusable framework that can easily be adapted to analyze new RAG components, such as retriever and reader models, as they evolve. We release our full dataset and code, aiming to provide the community with a deeper understanding of the nuanced interplay between context quantity, quality, and model architecture in RAG systems.
2 The RAGGED Framework
2.1 Framework Overview
We first explain the three aspects we vary in our analysis, then explain the three research questions we can answer with our analysis. The three aspects we vary in our framework are:
RAG system components. For example, we vary the choice of the retriever (e.g., BM25, ColBERT), the reader family, (e.g., LLaMa2, FlanT5), and the max input token length.
The number of retrieved passages, denoted as . We vary from from 1 to 100, though find the most insightful variations in behavior occur before .
Data slices of data to examine. For example, we examine slices where the top- retrieved passages include gold passages and where they do not include gold passages.
By varying these, we analyze the three following aspects of RAG system behavior:
Effective Number of Context Passages (§5)
We vary the number of passages and the choice of reader to see how different model architectures and context limits affect the effective number of context passages a reader can process. Here, we evaluate all instances instead of specific subsets. While some models monotonically improve with more provided passages, others may have an early peak and have limited ability to utilize more passages.
Context Utilization Behaviors (§6)
We focus on varying the choice of reader and the slice of instances to examine how different readers use context when the context quality varies. We examine the impact of context quality by analyzing slices of data where the top-k retrieved passages include gold passages and when they do not. As a result of this analysis, one can better understand how sensitive a reader is to positive context and how robust it is against incorrectly retrieved passages. We also investigate when models benefit from context and when models may be harmed by context.
Influence of Retriever Quality (§7)
We focus on varying the retriever and data domain, and observing how well retrievers can perform based on the nature of the questions and how sensitive readers are to the quality of the retrieved passages. As reader models may have varied sensitivity to retrieval quality, one could select more appropriate models given the question characteristics and retrieval performance.
2.2 Implementation Details
For all experiments, we use the following prompt:
Instruction: Give simple short one phrase answers for the questions based on the context Context: [passage, passage, , passage] Question: [the question of the current example] Answer: .
We truncate the Context to make sure the latter question is still included in the prompt. While performing generation for with top- passages with requires demanding computation, we sample to represent the general trend. More model implementation details can be found in §A.
3 Experimenting with RAG Systems
Below, we explain the representative retrievers and readers we analyze in our experiments.
3.1 Retriever
For retrieval, we employ two contrasting approaches: (1) sparse retrieval, based on lexical information, and (2) dense retrieval, based on neural models.
BM25
BM25, a probabilistic retrieval model formulated by Robertson et al. (2009), leverages TF-IDF (Chen et al., 2017b) principles to estimate passage relevance via term weighting and passage length normalization. BM25 uses term-matching and hence experiences fewer out-of-domain failures when facing special-domain vocabulary, such as legal or medical texts.
ColBERT
One of the best-performing neural-based retrievers is ColBERT (Khattab and Zaharia, 2020), i.e., contextualized late interaction over BERT. The transformer-based, contextualized embedding of ColBERT makes it more proficient than BM25 at identifying semantic similarities between queries and passages beyond lexical matching.
3.2 Reader
We compare four top-performing, open-source reader models of varied architectures: encoder-decoder models from the Flan family, and decoder-only models from the LLaMa2 family.
Flan Models
The Flan models are encoder-decoder models. We use the FlanT5-XXL (Chung et al., 2022) with 11B parameters and Flan-UL2 (Tay et al., 2023) with 20B parameters, both with a context length of 2 tokens. FlanT5-XXL is an instruction-tuned variant of the T5 model (Raffel et al., 2023). Flan-UL2 (Tay et al., 2023) is an upgraded T5-based model that is trained with Unifying Language Learning Paradigm, a pertaining process that uses a mixture-of-denoisers and mode switching to improve the model’s adaptability to different scenarios.
LLaMa Models
The Llama2 models are decoder models. For LLaMa models (Touvron et al., 2023a), we adopt the latest LLaMa models (Touvron et al., 2023b) in 7B and 70B parameters, both having a context length of 4k tokens. The key feature is that LLaMa models are trained with reinforcement learning human feedback (RLHF).
4 Datasets and Evaluation Metrics
4.1 Dataset
We adopt three DBQA datasets from various domains (Wikipedia, biomedical) and of various types (single-hop, multi-hop, list, yes/no).
Natural Questions
We choose the Natural Questions (NQ) dataset (Kwiatkowski et al., 2019) to examine model performance on the most generic open-domain, single-hop questions. NQ questions are real user-search queries on Google. We adopt the KILT version (Petroni et al., 2021) of the dataset, which for each example, provides at least one gold relevant passage and one short answer.
HotpotQA
We choose HotpotQA (Yang et al., 2018) which provides challenging multi-hop questions. Each question requires reasoning over at least two passages to answer. While maintaining the same Wikipedia domain with the NQ dataset, HotpotQA enables comparisons of model reasoning skills over multiple evidence pieces.
BioASQ
We choose BioASQ’s Task 11B (Krithara et al., 2023) with biomedical questions as a representative of special-domain questions. Our evaluation dataset is a compilation of the BioASQ Task 11B training and golden enriched set. BioASQ also presents challenging question types, such as lists and yes/no questions.
For NQ and HotpotQA datasets in the open domain, we retrieve passages from the Wikipedia corpus provided by the KILT benchmark (Petroni et al., 2021), with a total of 111M passages. For BioASQ, we use the PubMed Annual Baseline Repository for 2023 (of Medicine, 2023), with a total of 58M passages, which are either titles or passages of PubMed papers.
4.2 Metrics
Below we describe the metric definition for each instance. Overall performance on a dataset is computed by averaging over all instances.
4.2.1 Retrieval
Following Petroni et al. (2021), we evaluate retrieval performance using the recall@k metric.
Recall@k
For a given example with a query and ground-truth passage(s), recall@k measures the ratio of ground-truth passages among the top- retrieved passages. This measure can readily support retrieval for multi-hop questions (i.e., in HotpotQA) as well, where instead of retrieving any one of the gold passages, all passages along the reasoning path are necessary.
4.2.2 Reader Metrics
We evaluate the reader predictions using exact match and unigram F metrics.
Exact Match
Exact match (EM) measures if the model-generated output is the same as at least one of the ground-truth answers (Richardson et al., 2013). This metric requires models to be precise, yet may be overly strict when models produce verbose answers. Given the extractive nature of the NQ dataset (i.e., answers are spans in supporting passages), we use EM to evaluate NQ.
Unigram F
Unigram F is a well-adopted metric for QA, which quantifies the overlap of unigrams in the generated and reference texts, computed through the harmonic mean of precision and recall at the unigram level (Petroni et al., 2021). For each query, we compute the F score of the reader output against all possible gold answers and report the highest score. This is suitable if we do not need the generated answer to match the exact formatting and wording of the ground truth answer. For example, “five seasons” and “5 seasons” would fail EM but can get a partial credit of 0.5 by F. We evaluate HotpotQA and BioASQ with F metric.
5 Are More Contexts Always Better?
In this section, we study how models perform with various numbers of passages (§5.2), and if these results relate to context limits (§5.3).
5.1 RAGGED Analysis Guidelines
What to Vary:
Experiment with different numbers of context passages provided to the reader model. Start from a small number first since many reader models’ performance peaks before their context limits are reached.
Behaviors to Expect:
Expect a non-linear relationship between the number of contexts and model performance. Initially, reader performance may improve with more contexts due to the increased availability of “signal”, or supporting information. However, as the number of context passages increases, the amount of “noise”, or irrelevant, information also increases. As a result, reader performance may plateau or even degrade.
Model Behavior Implications:
Fitting as many contexts as the reader’s context limit can hold does not always guarantee optimal downstream performance. Using these experiments, one can better understand a reader’s ability to sift for “signal” among “noise”, and thereby find the optimal range of context passages for your model. Staying within this range maximizes performance without incurring unnecessary computational costs.
5.2 Number of Provided Contexts
We analyze the effect of the number of retrieved passages on reader model performance by providing the top- passages retrieved by ColBERT and evaluating the EM scores of answers generated by four reader models.
Results in Figure 3 reveal substantially different patterns between Flan and LLaMa models. On the NQ dataset, EM scores of both LLaMa models peak early at before continuing to decline steadily. In contrast, Flan models peak later around , and only level off instead of declining. When conditioning on a total of 30 passages, LLaMa 7B and LLaMa 70B underperform Flan models by about 17 to 33 point increase in exact match.
Similarly on multi-hop questions, Flan models can still benefit from more than 10 passages, whereas LLaMa can only benefit from 1 or 2 passages. In contrast, on the BioASQ dataset, both models have a small peak around 2 passages and roughly maintain the results afterwards.
These findings imply that encoder-decoder models can more effectively process tens of passages. In contrast, decoder-only models can use at most 2–3 passages in all cases. This observation underscores the necessity to carefully select the number of provided passages for different models to optimize their performance.
5.3 Context Limit vs. Early Truncation
Despite LLaMa models having twice the context length of Flan models (4 versus 2), they do not achieve superior performance when there is a large number of contexts. Nonetheless, this evidence alone can not allow us to claim that LLaMa models are worse than Flan models at processing long contexts. LLaMa’s longer context window could be a confounding factor that exposes the model to more (noisy) passages. To make a fair comparison, we truncate LLaMa model inputs by 2 tokens as we do to the Flan models, then evaluate LLaMa’s truncated-context performance.
As shown by the dashed lines in Figure 3, early truncation can only prevent LLaMa from further degrading after 15 passages on Wikipedia-domain questions but it does not improve results with smaller . The best-performing LLaMa (70B, truncated) still scores lower than the smallest Flan model, despite an almost 7 times difference in their number of parameters (70B vs. 11B).
This outcome challenges the assumption that models trained with extended context limits are inherently better at processing more contexts. Contrary to expectations, our results show that the short-window encoder-decoder models are more efficient in leveraging long contexts.
6 Context Utilization Habits
We study how readers utilize contexts, by examining how capable models are at answering questions without context, but instead simply from their pre-trained knowledge and generalization capabilities (§6.2). We then study how the readers incorporate test-time contexts with varied qualities (§6.3, §6.4).
6.1 RAGGED Analysis Guidelines
What to Vary:
Analyze the reader model’s performance using different slices of instances that represent different qualities of retrieved contexts.
The first slice is where the retrieved context includes at least one gold, human-annotated passage. In this scenario, there is sufficient information in the top-k passages to answer the question. The alternative reader context settings to compare with are when there are a) only the gold passages in the top-k retrieved passages and b) no context. Comparing with a) shows how good the reader is at ignoring noise and how close it is to using only signal and no noise. Comparing with b) shows if the reader is prone to performing worse
The second slice of instances is where none of the top-k retrieved contexts include gold passages. In this scenario, all the passages are noise. The alternative reader context setting to compare with is when there is no context.
Behaviors to Expect:
Models may show varying degrees of sensitivity to the quality of context. Reader models provided with only gold passages often act as an upper bound for the reader models’ performance with top-k passages. Although one might expect reader models provided with no context passages to act as a lower bound for the reader models’ performance with top-k passages, that is not always the case. It depends on the reader’s ability to discern between potentially sufficient internal knowledge and irrelevant context knowledge.
Implications of Behaviors:
For practitioners, these comparisons would tell us how effective the reader model sifts for signal among noise. If the model performs very closely to how it would with only gold passages, then it is highly effective at using relevant information and ignoring irrelevant ones. Conversely, a significant drop in performance with lower-quality contexts suggests a reliance on high-quality retrieval and potential vulnerability to noise. This analysis can guide the development or selection of more robust models that can maintain performance despite the variability in context quality.
6.2 No-context Generalization from Pre-trained Knowledge
We evaluate how capable models are at answering questions directly from their parameters and without any context by evaluating reader test-time performance without any context. We denote this setup as no-ctx and compare it with the standard setting with top-1 passages from ColBERT.
| Model | NQ | HotpotQA | BioASQ | |||
| no-ctx | top-1 | no-ctx | top-1 | no-ctx | top-1 | |
| Flan-T5 XXL | 16.7 | 33.8 | 27.4 | 43.5 | 18.2 | 40.0 |
| Flan-UL2 | 23.7 | 35.0 | 29.6 | 43.4 | 19.2 | 30.2 |
| LLaMa2 7B | 21.5 | 29.2 | 23.5 | 38.2 | 21.1 | 32.8 |
| LLaMa2 70B | 34.1 | 35.6 | 32.5 | 46.4 | 26.0 | 35.9 |
As shown in Table 1, under the no-context setting, LLaMa 70B outperforms Flan models by a large margin, suggesting they have at least encoded more knowledge from pretraining or are better at generalizing from memorized knowledge. While LLaMa has potentially memorized more relevant information during training, this knowledge may be more likely to cause conflicts with passages provided in context at test time. Next, we examine two particular slices of examples and study how reader models resolve conflict with irrelevant passages when gold (human-annotated) passages are present (§6.3) or absent (§6.4).
6.3 With Ground-Truth Passages
We study how readers handle context with both positive (gold) passages and negative (non-gold) passages. Specifically, for each given , we first find the slice of examples where the retriever returns at least one gold passage in top- candidates. With this slice of examples, we compare the results when providing (1) top- retrieved passages, (2) only the gold passage(s) within the top-k passages, and (3) no passages at all.
Figure 4 shows that LLaMa models are more severely affected by noisy context, as their scores decrease more sharply as increases. Notably, around and , LLaMa 7B and 70B models’ performance deteriorates to a level below that of their no-context baseline. In comparison, Flan models consistently outperform their no-context counterparts by a large margin.
Special Questions and Domains
We perform similar analyses on HotpotQA and BioASQ datasets. While Flan models exhibit trends akin to their open-domain behaviors, we observe substantial behavioral shifts on LLaMa models (Figure 5). Unlike the severe degradation on NQ questions, LLaMa consistently outperforms the no-context baseline. Additional special-domain contexts seem to induce less conflict with LLaMa’s parametric knowledge than additional contexts do in NQ potentially because they provide new, useful information that the model cannot derive from its pertaining memory alone.
6.4 Response to Negative passages
We now study the setting where no passages are explicitly useful to answer the questions. At each , we select the slice of examples where no gold passages exist among the top- retrieved candidates. We report model results when providing (1) these negative contexts or (2) no contexts at all.
As shown in Figure 6, compared to the no-ctx baseline, LLaMa performance degrades when contextualizing more negative passages (i.e., as increases). On the other hand, Flan models can benefit from negative contexts, consistently outperforming no-ctx baselines across all . Surprisingly, Flan models perform better with negative passages than with contexts at all. One explanation is that the negative passages may still be partially helpful even though they do not contain the exact correct answer since they may still contain some keywords or related topics that provide hints to the reader models. While LLaMa models fail to identify such subtle information, Flan models are more sensitive to related knowledge and capture these cues effectively.
Special Questions
We perform similar analyses on HotpotQA and BioASQ datasets.333For multi-hop questions, we select examples retrieved with all gold passages within the top- passages since all passages are necessary to answer the question. On HotpotQA, Flan models obtain less gain using negative passages and levels off above the no-ctx baseline (Figure 7). In contrast, LLaMa can use negative passages more effectively on special questions, outperforming no-ctx for most ’s on HotpotQA and on BioASQ.
Comparing models of different sizes, smaller models benefit more from negative contexts while larger models suffer. Between the two LLaMa models, while LLaMa 70B significantly regresses beyond its no-context baseline, LLaMa 7B performs comparably with its no-context baseline. Between the two Flan models, while FlanT5 benefits a lot from negative passages and outperforms no-ctx by a large margin, Flan-UL2 becomes sensitive to negative information and deteriorates below its no-ctx baseline across nearly all k values.
Case Study: What Negative Contexts Help?
If a negative context is related but insufficient for answering the question, it can still help improve the reader model’s performance. For instances where FlanT5 answers correctly with the top-5, negative contexts but incorrectly with no contexts, 43% of instances have top-k passages that include at least one passage from the Wikipedia page that the gold passages are from. Even though none of these negative passages include the exact correct answer, they are topically related enough that FlanT5 can bridge the knowledge gap.
Nonetheless, not all negative passages from the gold page are helpful. Among instances where top-5 underperforms no-ctx, 28% include at least one passage from the ground-truth Wikipedia page.
7 Comparing Different Retrievers
We compare retrievers of different quality (§7.2) and study their influence on the reader model’s downstream performance (§7.3).
7.1 Ragged Analysis Guidelines
What to Vary:
Use different retrieval algorithms (varying in architecture, computational cost, and storage cost). Here, we compare traditional, lexical and dense, neural retrievers.
Behaviors to Expect:
Retriever performance highly depends on the domain and complexity of the question (number of hops). However, interestingly, good retriever performance does not always guarantee good reader performance. The reader performance also depends on the reader’s sensitivity to retriever quality, with some models showing robustness to suboptimal retrieval.
Implications of Behaviors:
This aspect of the analysis informs the optimal choice of retriever for your RAG system and how much optimizing for that choice helps the reader’s performance. A model that performs well with various retrievers offers flexibility and might be more versatile in real-world applications. In contrast, a model highly sensitive to retrieval quality might require a state-of-the-art retriever to achieve optimal performance.
7.2 Evaluating Different Retrievers
We evaluate BM25 and ColBERT’s retrieval performance by computing recall@k for all k from 1 to 50 (Table 2). In general, ColBERT performs better than BM25 on questions about Wikipedia knowledge (NQ and HotpotQA). However, ColBERT only offers a negligible advantage over BM25 on the special-domain BioASQ, yet at a much higher computation cost.
Averaged over ranging from 1 to 50, ColBERT offers a 50-point gain over BM25 for NQ, but only offers a point gain over BM25 for BioASQ for paragraph recall. When applying to the out-of-domain questions, ColBERT does not gain many advantages from its neural, contextual embeddings.
| Model | Recall@k | |||||
| 1 | 2 | 5 | 10 | 20 | 50 | |
| NQ | ||||||
| BM25 | 0.7 | 1.1 | 2.5 | 4.1 | 6.0 | 7.5 |
| 10.3 | 16.3 | 27.8 | 36.8 | 47.7 | 53.2 | |
| ColBERT | 12.3 | 18.0 | 25.7 | 32.1 | 38.1 | 41.8 |
| 27.2 | 38.8 | 54.4 | 65.0 | 72.9 | 77.2 | |
| HotpotQA | ||||||
| BM25 | 0.2 | 0.4 | 1.0 | 1.6 | 2.4 | 3.0 |
| 23.3 | 31.2 | 42.7 | 52.1 | 59.1 | 62.8 | |
| ColBERT | 31.1 | 40.1 | 49.9 | 56.2 | 61.9 | 64.9 |
| 34.2 | 44.7 | 56.3 | 63.6 | 69.9 | 73.1 | |
| BioASQ | ||||||
| BM25 | 8.8 | 12.9 | 19.6 | 25.8 | 33.3 | 37.8 |
| 12.4 | 16.4 | 23.9 | 30.6 | 38.7 | 43.6 | |
| ColBERT | 8.8 | 13.5 | 20.7 | 27.1 | 34.3 | 38.6 |
| 14.2 | 18.2 | 25.6 | 32.2 | 39.8 | 44.2 | |
7.3 Generation with Different Retrievers
More importantly, we are interested in the question: how much does the choice of retriever affect downstream performance? We find that the nature of the task (e.g., single-hop vs. multi-hop, domain specificity) can significantly affect the choice between neural and sparse retrievers (Figure 8). While neural retrievers may require more resources, their great advantage in single-hop, Wikipedia-based questions and small advantage in specialized domains may still justify the investment. In contrast, neural retrieves are less beneficial for multi-hop questions, even for the Wikipedia domain.
Open Domain (Wikipedia)
For open-domain single-hop questions, using ColBERT offers substantial improvements over using BM25. At , ColBERT helps Flan models achieve a significant 16–18 points EM improvement and LLaMa2 models a more modest 4–6 point increase. This suggests there is value in using high-quality neural retrieval for open-domain, single-hop questions.
Special Question Type: Multi-Hop
For multi-hop questions, despite the stark retrieval performance gap between BM25 and ColBERT, the impact on reader performance is minimal. At , Flan models improve by 3–4 in F when paired with ColBERT over BM25, while LLaMa2 models improve by 2–3 in F. Both models show marginal gains with ColBERT, suggesting that multi-hop questions challenge reader models more in reasoning than context utilization.
While BM25 consistently gets recall scores 10 points lower than ColBERT, it is surprising that the reader performs comparably to using ColBERT-retrieved passages. This could be explained by how BM25 still has a high Wikipedia page recall, indicating its capability of finding relevant but inexact information that readers may still benefit from.
Special Biomedical Domain
ColBERT and BM25 perform similarly in the biomedical domain, so we might also expect the downstream performances to be similar. However, there is still a difference, albeit a small one, between using ColBERT and BM25. For , Flan models achieve a 2–5 point improvement when paired with ColBERT over being paired with BM25, while LLaMa2 models exhibit a much smaller, point improvement. Although there are only slight differences in retrieval quality, the specialized nature of the domain can amplify the impact on reader performance. As a result, there is still a discernible, though modest, preference for ColBERT over BM25 in accessing quality context.
8 Related Work
Context Limit and Processing Capacity
LMs with longer context windows are applicable across various knowledge-intensive generation tasks (Beltagy et al., 2020; Bertsch et al., 2023; Su et al., 2024). However, it is unclear how performant these models are in processing long contexts. Liu et al. (2023) study if LMs can be sensitive to the position of useful content within a long context, and struggle when it is in the middle. Moreover, Xu et al. (2024) show that an LM with a smaller context window (4) using RAG performs comparably with finetuning with a longer-window LM (16). Following this query, our work studies the effectiveness of LMs in utilizing long contexts, when they have different input capacities.
Domain Influence on Downstream Performance
It is crucial to know when LMs benefit from including retrieved passages in context. Mallen et al. (2023) find that retrieving contexts may be unnecessary and even detrimental when asking about common knowledge, but it benefits questions about rare knowledge. In contrast, we find that using RAG under the right configurations still offers significant downstream performance boosts even for common, Wikipedia-based questions.
Robustness to Noisy Contexts
Feeding in noisy contexts deteriorates LM performance. Asai et al. (2022) propose to select documents with high evidentiality scores. Wang et al. (2023) learn a filter model to remove the noisy sentences, and Berchansky et al. (2023) adopt a similar approach at the token level. Further, (Yu et al., 2023; Xu et al., 2023) use a neural summarizer model to aid the LM in identifying relevant information in long retrieved passages. Instead of reducing noise at test time, Yoran et al. (2023) train LMs to be robust to irrelevant content. Lastly, Chen et al. (2023) build an evaluation benchmark to test LMs’ noise robustness. Our work similarly studies LMs’ responses to noisy content but is more fine-grained with varied noise ratios.
9 Conclusion
We propose RAGGED, a framework designed to assist researchers and practitioners in making informed decisions about designing RAG systems, focusing on three key aspects: the number of contexts, the reader model, and the retriever model. We demonstrate the framework’s utility in deriving insights about RAG behaviors in response to varied context volumes, document quality, and question domains. We hope that our framework will be utilized by the community to deepen the understanding and customization of RAG systems.
Limitations
Although this study provides valuable insights into RAG systems, it has several limitations. First, the RAGGED framework, although comprehensive, focuses mainly on document-based question-answering tasks, which may not fully capture the nuances of other knowledge-intensive NLP tasks such as summarization, fact verification, and machine reading comprehension. Second, our experiments were conducted with a specific set of models and datasets, which may limit the generalizability of our findings to other models, languages, or domains not covered in this study. However, providing a comprehensive analysis is not the main motivation or contribution of the paper. We encourage readers to use our framework to evaluate other models and datasets and share the insights with the community.
Acknowledgements
Special thanks to Alex Cabrera, Alex Bäuerle, Jun Araki, Md Rizwan Parvez for providing Zeno support for analysis visualization. Our appreciation extends to Hao Zhu, Jacob Springer, and Vijay Viswanathan for providing feedback for our paper. This paper was supported in part by a gift from Bosch research.
Appendix A Implementation Details
Reader model
Specifically, when using FlanT5 and FlanUL2 readers, we use T5Tokenizer to truncate sequences to up to 2 tokens; when using LLaMa2 models, we apply the LlamaTokenizer and truncate sequences by 4 tokens. Subsequently, we incorporate a concise question-and-answer format that segments the query using "Question:" and cues the model’s response with "Answer:", ensuring precise and targeted answers.
For our reader decoding strategy, we used greedy decoding with a beam size of 1 and temperature of 1, selecting the most probabble next word at each step without sampling. The output generation was configured to produce responses with 10 and 256 tokens to examine the effect of varying response lengths on model performance. The experiments were conducted on NVIDIA A6000 GPUs, supported by an environment with 60GB RAM. The average response time was 1.1s per query when processing with a batch size of 50.
Appendix B Dataset Details
All corpus and datasets use English.
The Medline Corpus is from of Medicine (2023) provided by the National Library of Medicine.
| Corpus | # of par | # of doc | Avg # of doc |
| Wikipedia | 111M | 5M | 18.9 |
| Medline | 58M | 34M | 1.7 |
For NQ and HotpotQA, we use KILT’s dev set versions of the datasets, allowed under the MIT License (Petroni et al., 2021). For BioASQ (Krithara et al., 2023), we use Task 11B, distributed under CC BY 2.5 license.
| Dataset | # of Queries |
| NQ | 2837 |
| HotpotQA | 5600 |
| BioASQ | 3837 |